Southwestern University (SWU), a large state college in Stephenville, Texas enrolls close to 20,000 students. It is the dominant force in the small city, with more students during fall and spring than permanent residents. A longtime football powerhouse, the University s a member of the Big Eleven conference and is usually in the top 20 in college football rankings. To bolster its chances of reaching the elusive and long-desired number-one ranking, in 2013 the University hired the legendary Billy Bob Dillon as its head coach. Although the number-one ranking remained out of reach, attendance at the six Saturday home games each year increased. Prior to Dillon’s arrival, attendance generally averaged 30,000–35,000. Season ticket sales bumped up by 15,000 just with the announcement of the new coach’s arrival. Stephenville and the University were ready to move to the big time! With the growth in attendance came more fame, the need for a bigger stadium, and more complaints about seating, parking, long lines, and concession stand prices. The University’s president was concerned not only about the cost of expanding the existing stadium versus building a new stadium but also about the ancillary activities. He wanted to be sure that these various support activities generated revenue adequate to pay for themselves. Consequently, he wanted the parking lots, game programs, and food service to all be handled as profit centers. At a recent meeting discussing the new stadium, the University’s president told the stadium manager to develop a break-even chart and related data for each of the centers. He instructed the stadium manager to have the food service area break-even report ready for the next meeting. After discussion with other facility managers and his subordinates, the stadium manager developed the following table showing the suggested selling prices, his estimate of variable costs, and his estimate of the percentage of the total revenues that would be expected for each of the items based on historical sales data. Item Selling Price/Unit $ Variable Cost/Unit $ Percent Revenue % Soft drink 1.00 0.45 15% Coffee 3.00 0.60 30% Hot dogs 3.50 0.85 25% Burgers 4.50 1.50 18% Misc. Snacks 2.00 0.90 12% The stadium manager’s fixed costs are interesting. He estimated that the prorated portion of the stadium cost would be as follows: salaries for food services at $360,000 ($60,000 for each of the six home games); 2,400 square feet of stadium space at $ 8 per square foot per game; and six people per booth in each of the six booths for 5 hours at $15 an hour. These fixed costs will be proportionately allocated to each of the products based on the percentages provided in the table. For example, the revenue from soft drinks would be expected to cover 20 % of the total fixed costs. The stadium manager wants to be sure that he has a number of things for the University President: (1) the total fixed cost that must be covered at each of the games; (2) the portion of the fixed cost al- located to each of the items; (3) what his unit sales would be at breakeven for each item—that is, what sales of soft drinks, coffee, hot dogs, hamburgers, and snacks are necessary to cover the portion of the fixed cost allocated to each of these items; (4) what the dollar sales for each of these would be at these break-even points; and (5) realistic sales estimates per attendee for attendance of 60,000 and 35,000. (In other words, he wants to know how many dollars each attendee is spending on food at his projected break-even sales at present and if attendance grows to 60,000.) He felt this last piece of information would be helpful to understand how realistic the assumptions of his model are, and this information could be compared with similar figures from previous seasons. Discussion Questions Calculate a. The total fixed cost per game and the allocated cost to each food item b. Calculate the Break-Even-Point (BEP) for each of these items c. Determine the total sales for each item that is required to each BEP d. Write your conclusion whether we break even or not if the attendance is 35000 people or 60000 people. e. Prepare a brief report for the University President that covers the items noted.
Answer #1

report to the president of the unviersity

1) total fixed cost covered in each game

food service 60000
stadium space 19200 =(2400*8)
booth staff wage 450 =(6*5*15)
total fixed cost per game 79650

allocated cost for each food item in each game

assumin that proportion is based on the total revenue percentage

soft drink = 79650*15% = 11947.5 / for six game =71685

coffee = 79650*30%= 23895 / for six game =143370

hot dogs = 79650*25%= 19912.5 / for six game =119475

burger = 79650*18%= 14337 / for six game = 86022

snacks = 79650*12%= 9558 / for six games = 57348

2) units of each items to cover the break even point

contribution of each item

item contribution
soft drink 0.55
coffee 2.4
hot dogs 2.65
burger 3.00
misc snacks 1.1

weighted average break even point (units)

=(477900)/(0.55*15%+2.4*30%+2.65*25%+3*18%+1.1*12%)= 223632 units

each item units

soft drink =223632*15%=33544.8 units

coffee =223632*30% = 67089.6units

hot dods = 223632*25% =55908

burger = 223632*18 = 40253.8

misc snck = 223632*12% =26835.8

sales of each item required to cover the break even

soft drink =223632*15%=33544.8 units * $1 =33544.8

coffee =223632*30% = 67089.6units * $3 =201268.8

hot dods = 223632*25% =55908 * $3.5 = 195678

burger = 223632*18 = 40253.8 * 4.5 = 181142.1

misc snck = 223632*12% =26835.8 * 2 = 53671.6

total break even sales = =665305.3

total sales if attendance is 35000

soft drink 35000*15%*$1= 5250
coffee 35000*30%*$3= 31500
hot dogs 35000*25%*3.5= 30625
burger 35000*18%*4.5= 28350
snacks 35000*12%*2= 8400
total sales 104125

total sales if attendance is 60000

soft drink 60000*15%*1= 9000
coffee 60000*30%*3= 54000
hot dogs 60000*25%*3.5= 52500
burger 60000*18%*4.5= 48600
snacks 60000*12%*2 = 14400  
total sales = 178500

conclusion

in both estimated attendances university will not acheive it break even sales